### v1

# From Seedling to Harvest: The GrowingSoy Dataset for Weed Detection in Soy Crops via Instance Segmentation

Abstract
Deep learning, particularly Convolutional Neural Networks (CNNs), has gained significant attention for its effectiveness in computer vision, especially in agricultural tasks. Recent advancements in instance segmentation have improved image classification accuracy. In this work, we introduce a comprehensive dataset for training neural networks to detect weeds and soy plants through instance segmentation. Our dataset covers various stages of soy growth, offering a chronological perspective on weed invasion’s impact, with 1,000 meticulously annotated images. We also provide 6 state of the art models, trained in this dataset, that can understand and detect soy and weed in every stage of the plantation process. By using this dataset for weed and soy segmentation, we achieved a segmentation average precision of 79.1% and an average recall of 69.2% across all plant classes, with the YOLOv8X model. Moreover, the YOLOv8M model attained 78.7% mean average precision (mAp-50) in caruru weed segmentation, 69.7% in grassy weed segmentation, and 90.1% in soy plant segmentation. [^1] Video of the experiments is available at: https://youtu.be/Ir0vwtAk8Tc [^2] GitHub Repository is available at: https://github.com/raulsteinmetz/soy-segmentation [^3] SoyGrowing dataset is available at: https://github.com/raulsteinmetz/soy-segmentation-ds
Index Terms:
Weed Detection, Instance Segmentation, Temporal Perspective Dataset, Soy
I Introduction
Figure 1: Some examples of instance segmentation on the GrowingSoy dataset. Arrows highlight the transitions between temporal states, encompassing all growth stages of the soy plant.
Effective crop management is essential for maximizing yields, especially in environments prone to issues like weed infestations, such as soybean crops. Despite efforts to control weeds through methods like chemical applications, the constant threat of unexpected weed invasions looms large, endangering crop quality and yield. Thus, vigilance and responsive agricultural practices remain paramount.
Soybeans occupy a pivotal role on the world stage, being the second most produced vegetable oil and serving as a major source of high-quality protein for both animal feed and vegetable oil production, while also delivering essential protein for human consumption [1]. It is of uttermost importance to use management techniques that allow to identify and remove weeds as early as possible, increasing crop production.
In recent years, rapid advancements in hardware and software technologies have propelled neural networks into the spotlight as powerful tools capable of automating crucial agricultural processes [2, 3]. These innovations encompass autonomous robot control [4, 5] and sophisticated computer vision algorithms, such as Convolutional Neural Networks (CNN) for image classification [6]. These technologies enable automated disaster detection while simultaneously conducting a comprehensive analysis of plantation status and key metrics [7].
One groundbreaking development in computer vision is instance semantic segmentation [8, 9], a technique that partitions unfamiliar images into distinct components and entities with pixel-level precision, revolutionizing object recognition and delineation. This advancement holds immense promise for agricultural computer vision, where precise identification and understanding of individual components within a crop field are crucial for effective crop management. However, the lack of data properly labeled and processed for segmentation is yet a bottleneck for the development of advanced applied solutions.
The prevailing issue in crop datasets lies in the lack of segmentation labels, as the majority remain unlabeled for this purpose. Notably, datasets such as DeepWeeds [10], Weed25 [11], and the Soybean images dataset  [12] are exclusively annotated for image classification, overlooking the crucial segmentation aspect. Furthermore, a significant portion of these datasets consists of images with low resolution and inadequate illumination.
In contrast, our dataset address these limitations by providing a substantial number of high-quality images depicting soy and weeds. What sets our dataset apart is the inclusion of segmentation labels for all images, enhancing their uniqueness. This not only contributes to a more comprehensive dataset but also fills a crucial gap in the availability of well-annotated crop datasets for segmentation tasks.
Hence, this research aims to contribute with the context presented through the following contributions:
  • We present a novel dataset with 1,000 meticulously annotated and high-resolution images for training instance segmentation neural networks, focusing on weed-infested soy plantations. The dataset spans the entire soy growth process, from initial stages to maturity with pervasive weed presence.
  • Additionally, we provide six state-of-the-art neural network models and perform an in-depth comparison among them, leveraging this temporal dataset to achieve impressive mean average precision in weed and soy segmentation across all soy plantation stages.
The YOLOv5 [13] and YOLOv8 [14] models trained using our dataset, achieved promising performance. Notably, the YOLOv8m model achieved a segmentation mean average precision of 78.9% for caruru weed, 62.8% for grassy weed, and 87.7% for soy plants. The YOLOv8m model also attained segmentation map-50 scores of 78.7% for caruru weed, 69.6% for grassy weed, and 90.1% for soy plants. Additionally, the YOLOv8x model attained 0.791 precision score between all classes and 0.692 recall score between all classes. The instance segmentation precision of the model trained using our dataset is depicted in Fig. 1.
The structure of this paper is as follows: Section II offers an extensive literature review, delving into previous research in computer vision for crop monitoring and instance segmentation. In Section III, we outline our methodology for creating the dataset. In Section IV, we depict the process of selecting and training the neural networks involved in the benchmark study. Section V presents the results obtained using the neural networks on our dataset. Section VI presents an analysis of the performance in images that contain both soy and weed, in seedling and harvest plantation stages. Finally, Section VII summarizes our conclusions drawn from this study, highlighting the significance of our findings for agricultural management and the potential for further advancements in instance segmentation for crop monitoring.
II Related Work
Over the years, continuous advancements in artificial intelligence have enhanced the widespread use of CNNs in precision agriculture and the use of instance segmentation for computer vision. In this session, we will explore some works that obtained success employing this technology.
In the field of plant disease detection, leveraging the use of a deep CNN, Francis and Deisy [15] attained an 87% precision rate in the detection and classification of diseases in apple and tomato leaf images. They proposed a methodology consistent of an input image of 64x64 pixels and a smaller amount of parameters when compared with the other models that can be used for this problem. Similarly, Bedi and Gole [16] developed a hybrid AI model that combines a convolutional autoencoder and a CNN to identify the presence of Bacterial Spot disease in peach plants by analyzing images of their leaves, the model avoids immense quantities of parameters and also increases overall accuracy. Their main contribution was related to the small amount of parameteres (less then 10K parameters) reducing the amount of time needed for training and to perform the real time detection of possible plant deceases.
Weed detection through CNN’s has been widely adopted in the literature [17]. Ramirez et al. [18] employed the DeepLabv3 neural network architecture to identify weed presence within a beet agricultural field by analyzing high-resolution aerial imagery. For the latter, a comparison between three distinct model was done, concluding that the DeepLabv3 presented a better performance, besides being the one that needs the most computational resource. In a similar vein, Hu et al. [19] utilized the YOLOv4 architecture to identify 12 distinct types of weeds in a rice plantation. A comparative analysis with the YOLOv3 architecture was conducted, demonstrating enhancements in multiple metrics.
Several datasets containing images of both weeds and crops have been made available in the literature like the one presented by Sudars et al. [20], which consists of 1118 annotated images encompassing six different crop types and eight different weed types. Another notable dataset is the TobSet, presented by Alam et al. [21]. It is a collection of images featuring tobacco plants and weeds captured from local fields, encompassing diverse growth stages and varying lighting conditions. The dataset includes 7000 images of tobacco plants and 1,000 images of weeds.
In the domain of crop monitoring, another significant application of CNNs is in the detection of plant diseases. Medhi and Deb [22] introduced an image dataset featuring various banana plant varieties and the diseases associated with them, summing up to an extensive collection of over 8000 images for deep neural network training. In a similar vein, Moupojou et al. [23] introduced the FieldPlant dataset, consisting of 5,170 plant disease images acquired directly from plantations. These images were meticulously annotated at the leaf level by plant pathologists. The study demonstrates that FieldPlant surpasses current datasets in terms of performance in classification tasks, establishing itself as a more precise tool for disease detection in practical agricultural scenarios.
Leveraging the power of semantic segmentation, various computer vision tasks can have several improvements in accuracy. Yang et al. [24] conducted a comprehensive study comparing instance segmentation methods on established datasets, alongside an examination of practical applications of semantic segmentation in autonomous driving. Also using instance segmentation concepts, Zhou et al. [25] introduced an innovative approach, a combination of Swin Transformer and CNN, for tunnel lining crack detection using instance segmentation. This method achieved a mean intersection over union (mIoU) of 77.41% and a mean pixel accuracy of 84.42% on the custom datasets presented in this paper. Notably, these results outperform previous CNN-based and transformer-based semantic segmentation algorithms in terms of segmentation accuracy.
Semantic segmentation shows notable enhancements in various fields of crop management and monitoring. Anand et al. [26] introduced AgriSegNet, a deep learning framework designed for the automatic identification of farmland anomalies, including standing water and weed clusters, by employing multiscale attention-based semantic segmentation on UAV-captured images. This framework is tailored for IoT-assisted precision agriculture. Bosilj et al. [27] conducted a study examining the feasibility of transfer learning across different crop types, demonstrating the potential to reduce training time by up to 80% in semantic segmentation-based architectures.
To the best of our knowledge, no dataset in the literature provides RGB, high-quality images of soy plants along with human-annotated semantic segmentation labels, covering their growth stages from seedling to harvest. As well as encompassing the presence of weed invasions such as caruru and grassy weeds. The temporal factor in the dataset is especially important since you can track the plantation as well as the weeds invasion, for a better course of action regarding ways to deal with the weeds. Furthermore, this paper represents the pioneering effort to employ state-of-the-art YOLO architectures for accurate segmentation of both soy and weeds throughout all growth stages of these plants.
III The Dataset
Creating the soy instance segmentation dataset included recording videos, selecting images, and manually labeling for instance segmentation. This section illustrates these processes.
III-A Data Gathering
The data collection process involved recording multiple videos that captured various distinct stages of the soy crop development, spanning from the initial growth phases to more advanced stages. These videos also documented the presence of caruru and grassy weeds. The selected areas of the plantation where the videos were shot received different levels of chemical treatment, including both medium and chemical-free approaches. This recording took place at the Universidade Federal de Santa Maria, within a dedicated soy plantation for research, located at approximately 29° 43’ 42” S, 53° 45’ 24” W, an aerial image can be seen in Fig. 2.
Figure 2: Location where the videos were recorded.
The videos were captured using a GoPro Hero 12 camera with 4K resolution and a frame rate of 120 frames per second. To ensure comprehensive coverage, the camera was mounted on an ATV four-wheeled vehicle, which followed a consistent 14-meter path through the plantation at a steady pace of 2 kilometers per hour, recording the crop’s growth at various intervals. The ATV and camera setup can be seen in Fig. 3.
Figure 3: Setup used for the video recordings.
III-B Dataset Creation and Annotation
To curate a representative database of soy plantation stages, we systematically selected 1000 frames from recorded videos. This sequential extraction, with a fixed step across all videos, ensured comprehensive coverage of each stage in the soy plantation lifecycle. Through iterative experimentation, we determined the optimal number of images for model training, ranging from 250 to 1500 images. Interestingly, discernible model improvement plateaued after reaching 1000 images, affirming that this quantity sufficiently encapsulates the task’s complexity.
Subsequently, the selected frames underwent resizing to 640x640 pixels using cubic interpolation, aligning them with the specifications required for neural network models. To provide a glimpse into the dataset, a subset of these images is showcased in Fig. 4. This meticulous curation process not only ensures diversity in representation but also optimizes model performance for the designated task.
Figure 4: Image sample from our dataset. Different stages of the plantation can be easily visualized.
In the labeling process, instance segmentation was employed. This technique involved delineating and fitting polygons around all the objects, which in this case were the plants within the scenes. This detailed labeling approach allowed the neural networks to learn how to accurately segment each individual plant, enabling precise object recognition and classification within the dataset. The labeling was performed manually using the Roboflow Framework [28], which streamlines the creation of such labeled datasets for machine learning purposes [29].
IV Neural Network Models
In this section, we delve into an analytical exploration of two paramount frameworks in the realm of computer vision segmentation: YOLOv5 [13] and YOLOv8 [14], these architectures and their different size models (medium, large and x-large) were the ones chosen for our instance segmentation approach. These frameworks, underpinned by cutting-edge research and technological advancements, have significantly shaped the trajectory of real-time object detection and segmentation methodologies.
YOLOv5 is a significant iteration in the YOLO series, known for its real-time object detection capabilities. Developed by Ultralytics, YOLOv5 became a defining model in the computer vision industry. It operates by dividing an image into a grid and predicting bounding boxes and class probabilities for each grid cell. YOLOv5 uses anchor boxes, which are predefined shapes, to predict object shapes in an image. The model’s success is attributed to its flexible Pythonic structure, which allowed the community to introduce new modeling improvements and share them quickly.
YOLOv8 is the latest state-of-the-art model in the YOLO series, designed for object detection, image classification, and instance segmentation tasks. Developed by Ultralytics, the same team behind YOLOv5, YOLOv8 introduces numerous architectural and developer experience enhancements over its predecessor. One of the significant features of YOLOv8 is its anchor-free detection, which predicts the center of an object directly instead of using an offset from a known anchor box. This approach reduces the number of box predictions, speeding up the Non-Maximum Suppression (NMS) process. Additionally, YOLOv8 has introduced new convolutions, changing the stem’s first 6x6 conv to a 3x3 and introducing the C2f module. The model also incorporates online augmentations during training, such as mosaic augmentation.
The primary distinction between YOLOv5 and YOLOv8 is their approach to object detection. While YOLOv5 relies on anchor boxes, YOLOv8 operates as an anchor-free model, predicting object centers directly. This change addresses the challenges of anchor boxes, which might not represent the distribution of a custom dataset. Furthermore, YOLOv8 has introduced new convolutions and modules, such as the C2f module, which differentiates it from YOLOv5. Another notable difference is the training routine. YOLOv8 employs online augmentations, like mosaic augmentation, which stitches four images together, enhancing the model’s ability to recognize objects in various locations and conditions. In terms of accuracy, YOLOv8 has shown improvements over YOLOv5, especially when evaluated against benchmarks like COCO and Roboflow 100.
V Evaluation and Results
(a) YOLOv8 mAP increase.
(b) YOLOv8 class loss decrease.
(c) YOLOv8 segmentation loss decrease.
(d) YOLOv5 mAP increase.
(e) YOLOv5 class loss decrease.
(f) YOLOv5 segmentation loss decrease.
Figure 5: Performance metrics for YOLOv8 and YOLOv5 during training and validation over several epochs. The legends for each graph include variations of the model: ‘m’, ‘l’, and ‘x’, representing different model sizes or configurations, with ‘train’ indicating training data and ‘val’ indicating validation data.
The neural networks trained on the temporal dataset demonstrated significant precision, recall, and mean average pixel precision throughout all stages of the plantation. This section provides an overview of the network training process, presents the achieved metrics, and conducts a comparative analysis between the YOLO architectures.
The neural networks underwent training for 200 epochs, employing early stopping and a batch size of 8. The early stopping mechanism, with a patience hyperparameter set to 50 epochs, allowed for the termination of training when performance improvements became marginal. Pretraining the model on ImageNet [30] provided a solid foundation for feature extraction. The training process utilized a constant learning rate of 0.01, an Adam optimizer with a momentum of 0.937, and a weight decay of 0.0005. Notably, dropout was not applied, and the training mode operated in a deterministic manner, ensuring reproducibility and stability throughout the training process. This comprehensive configuration aimed to strike a balance between model complexity, convergence speed, and generalization performance.
The learning curves can be seen in Fig. 5. In Figures 5(b) and 5(e), the class loss curves for both the validation and training sets across all models are displayed. In Figures 5(c) and 5(f), the loss curve for segmentation can be seen. The loss curve for class is measured using Distribution Focal Loss (the focal loss helps to address class imbalance in object detection by assigning different weights to easy and hard examples), and for segmentation is Binary Cross Entropy with Logits (binary classification loss function that combines sigmoid transformation of logits with binary cross-entropy calculation, it’s used for stable and efficient training in binary classification tasks). These curves display a gradual and consistent decline, highlighting the models’ efficacy in comprehending the data throughout the training process. Additionally, the mean average precision on both the training and validation sets is presented in Figures 5(a) and 5(d), showcasing a steady and continuous upward trend, ultimately converging to yield outstanding results. The mAp-50 is calculated by averaging the precision values at the top 50 ranked predictions for a given task, providing a powerful performance metric for object detection.
Table I: Box mAP50 for all models and classes.
Architecture Caruru Weed Grassy Weed Soy Plant
YOLOv8m 0.789 0.628 0.887
YOLOv8l 0.759 0.593 0.880
YOLOv8x 0.758 0.613 0.870
YOLOv5m 0.744 0.623 0.866
YOLOv5l 0.772 0.607 0.856
YOLOv5x 0.720 0.606 0.860
In Table I, mAP-50 metrics for bounding boxes in the test set are shown. The YOLOv8m model achieved the highest scores, with 0.789 for caruru weed objects, 0.628 for grassy weed objects, and 0.887 for Soy objects. Notably, the YOLOv8 models outperformed the YOLOv5 models across the vast majority classes, indicating a significant improvement.
Table II presents the average bounding box precision and recall metrics for all classes across all models. YOLOv8x exhibited the highest average precision at 0.791, while YOLOv8m achieved the best recall at 0.733. Notably, YOLOv8 and YOLOv5 models demonstrated similar results in these metrics.
Table II: Average box precision and recall for all models and classes.
Architecture Precision (all classes) Recall (all classes)
YOLOv8m 0.767 0.737
YOLOv8l 0.729 0.719
YOLOv8x 0.772 0.693
YOLOv5m 0.773 0.728
YOLOv5l 0.770 0.738
YOLOv5x 0.746 0.730
The segmentation mAP-50 values can be seen in Table III, with YOLOv8m consistently outperforming other models across all classes, mirroring the box mAP-50 results. Specifically, YOLOv8m achieved a mAP of 0.787 for caruru and 0.696 for grassy weed, and an impressive 0.901 for Soy. Once more, the YOLOv8 models demonstrated superior performance compared to the YOLOv5 models.
Table III: Segmentation mAP50 for all models and classes.
Architecture Caruru Weed Grassy Weed Soy Plant
YOLOv8m 0.787 0.696 0.901
YOLOv8l 0.777 0.614 0.895
YOLOv8x 0.761 0.591 0.888
YOLOv5m 0.727 0.634 0.884
YOLOv5l 0.742 0.610 0.875
YOLOv5x 0.707 0.603 0.870
Precision and recall values for segmentation are provided at Table IV. YOLOv8x achieved the highest precision at 0.791, while YOLOv8m attained the best recall at 0.733. Notably, precision and recall values across all models were quite similar for this task.
To provide a more comprehensive illustration of model predictions on test images from the dataset, Fig. 6 presents sample predictions for YOLOv8x across the initial, medium, and advanced stages of soy growth. This includes precise segmentation and classification of soy, caruru, and grassy weeds, demonstrating the model’s accuracy in delineating these plants.
VI Analysis of performance on initial and late stages
In this section, the primary focus revolves around the neural network’s pivotal role in effectively detecting weed presence. Our analysis will be dedicated to evaluating the performance of the YOLOv8 models on images featuring a coexistence of both weed and soy elements, through box predictions rather then mask predictions. We meticulously curated two distinct sets of 25 images, each depicting varying degrees of weed presence. The initial group comprises images captured during the seedling and early stages of plantation, while the second group encapsulates images from the harvest stages. Our objective is to scrutinize and comprehend the YOLOv8 models’ efficacy in accurately identifying weed instances in these diverse agricultural scenarios, and if there is a several difference in performance from early to late stages.
Table IV: Average segmentation precision and recall for all models and classes.
Architecture Precision (all classes) Recall (all classes)
YOLOv8m 0.784 0.733
YOLOv8l 0.743 0.732
YOLOv8x 0.791 0.692
YOLOv5m 0.770 0.721
YOLOv5l 0.766 0.731
YOLOv5x 0.743 0.716
Figure 6: YOLOv8X model predictions over multiple stages of soy growth.
Considering the box mAP50 results outlined in Table V and Table VI, a noteworthy observation emerges. Despite YOLOv8x being notably surpassed by the more compact YOLOv8m model, a distinctive trend becomes evident in scenarios featuring both weed and soy elements, irrespective of whether the images capture the seedling or harvest stages. Notably, YOLOv8x exhibits a consistent trend of outperforming its counterparts.
Table V: Box mAP50 for all models and classes in early stages.
Architecture Caruru Weed Grassy Weed Soy Plant
YOLOv8m 0.995 0.915 0.936
YOLOv8l 0.995 0.838 0.930
YOLOv8x 0.995 0.981 0.946
Table VI: Box mAP50 for all models and classes in late stages.
Architecture Caruru Weed Grassy Weed Soy Plant
YOLOv8m 0.970 0.884 0.905
YOLOv8l 0.933 0.901 0.901
YOLOv8x 0.984 0.914 0.958
An additional intriguing facet, gleaned from the tabulated data, is the absence of significant disparities in model performance across distinct stages of plantation. This observation underscores the diversity of the dataset, affirming that the trained models possess the capability to comprehend and effectively navigate through a spectrum of agricultural scenarios.
Taking into consideration the box precision and recall metrics showcased in Table VII and Table VIII, it becomes apparent that no single model distinctly outperforms the others. All YOLOv8 models consistently deliver commendable results in accurately detecting both soy and weed in the images. Notably, a subtle trend emerges where precision and recall metrics exhibit a slight improvement in the later stages of plantation, although the magnitude of this enhancement is not particularly significant.
Table VII: Average box precision and recall for all models and classes in early stages.
Architecture Precision (all classes) Recall (all classes)
YOLOv8m 0.868 0.943
YOLOv8l 0.954 0.873
YOLOv8x 0.916 0.909
Table VIII: Average box precision and recall for all models and classes in late stages.
Architecture Precision (all classes) Recall (all classes)
YOLOv8m 0.885 0.841
YOLOv8l 0.858 0.864
YOLOv8x 0.879 0.933
In conclusion, this analysis not only underscores the similarity in performance of the networks across different stages of plantation, highlighting the dataset’s inherent diversity and the models’ robust generalization capabilities, but it also reveals a significant performance boost in images featuring the simultaneous presence of both soy and weed. This enhancement aligns seamlessly with the primary objective of the network, which is the segmentation and detection of soy and weed, reaffirming the effectiveness of the models in fulfilling their intended purpose.
VII Conclusion
This paper introduces a dataset containing 1000 images capturing the entire life cycle of soy crops, from seedling to harvest, while also documenting weed infestation over time. Each image provides a comprehensive view of the plantation and includes instance segmentation labeling. This dataset holds promising scientific value for applications related to weed detection, soy monitoring, and serves as a robust benchmark for neural network architectures due to its comprehensiveness and practicality.
The paper further includes analysis and comparison of state-of-the-art YOLO models using this dataset, contrasting the latest model iteration with the previous YOLOv5. The models trained on this dataset demonstrate exceptional precision in plant segmentation. The best results were achieved with YOLOv8M, 78.7% mean average precision (mAp-50) in caruru weed segmentation, 69.7% in grassy weed segmentation, and 90.1% in soy plant segmentation.
Looking ahead, potential future work could explore the extension of this dataset’s methodology for crop disease detection and yield prediction, capitalizing on the temporal aspect of the dataset, which offers valuable practical applications.
Acknowledgements
This research was funded by: Topsector TKI AgroFood under grant agreement LWV20242 and Ministry of Agriculture, Nature and Food Quality grant agreement BO-69-001-005 for “Smart technology for soybean production”.
The authors express their gratitude for the support and data contributions from various organizations, including Wageningen University & Research, the Aquarius Project at Universidade Federal de Santa Maria, the Advanced Farm 360 project, Rometron, and Stara. Special thanks are extended to Ard Nieuwenhuizen and Telmo Jorge Carneiro Amado for their invaluable support throughout the research process.